### v1

@@ Abstract
The rapid growth of large language model (LLM) services imposes increasing demands on distributed GPU inference infrastructure. Most existing scheduling systems rely on the current system state to make decisions, without considering how task demand and resource availability evolve over time. This lack of temporal awareness leads to inefficient GPU utilization, high task migration overhead, and poor system responsiveness under dynamic workloads. In this work, we identify the fundamental limitations of these instantaneous-state-only scheduling approaches and propose Temporal Optimal Resource scheduling via Two-layer Architecture (TORTA). TORTA introduces a spatiotemporal scheduling framework that captures both long-term workload patterns and short-term execution constraints. It adopts a two-layer design: a macro-level scheduler leverages reinforcement learning and optimal transport to coordinate inter-region task distribution, while a micro-level allocator refines task-to-server assignments within each region to reduce latency and switching costs. Experimental results across multiple network topologies show that TORTA reduces average inference response time by up to 15%, improves load balance by approximately 4-5%, and cuts total operational cost by 10-20% compared to state-of-the-art baseline methods.
keywords:
[…]
1 Introduction
@@
Accelerating GPU inference can be divided into two categories. The first is local acceleration, which involves optimizing the GPU’s operational mode. This includes techniques such as optimizing checkpoint loading, introducing real-time migration and intelligent scheduling strategies, or employing model parallelism and statistical multiplexing to reduce unnecessary overhead [25, 8, 22, 24, 55]. The second category is distributed resource scheduling, which aims to improve GPU resource utilization and prevent GPU server idleness [50]. Some approaches focus on allocating suitable GPU resources to different tasks to achieve the fastest inference, often with additional optimization objectives such as minimizing communication costs and power consumption [50, 41, 16].
@@
The complexity of this scheduling problem is often very high. Finding an optimal scheduling solution typically requires significant computational time and increases dramatically with network scale expansion, severely impacting task completion time. Existing approaches usually seek approximately optimal solutions within acceptable time constraints, and these methods have achieved good results to a certain extent. For example, MLServerless [8] employs an offline-plus-online strategy, where the offline phase invests substantial effort in learning solution strategies, enabling the system to quickly derive good solutions during the online phase to achieve optimal inference speed. Other approaches [16, 41] model this as a Mixed Integer Linear Programming (MILP) or Integer Quadratic Programming (IQP) problem and accelerate solution speed through solution space pruning. However, most of these approaches are "Reactive", making immediate decisions based on currently observed states (such as cluster interference levels, resource utilization rates, and energy prices), essentially finding optimal solutions within individual time slices. This paradigm inherently ignores the temporal perspective and continuity considerations in decision-making.
@@
Given these challenges, our approach addresses GPU resource allocation by integrating reinforcement learning with an optimal transport [42] perspective, particularly incorporating the temporal dimension. Unlike traditional methods, our strategy divides the allocation process into Two-layer Architecture: macro and micro. In the first layer (macro), we assign GPU inference tasks to nodes in each region, primarily to mitigate the imbalance between GPU resources and task distribution. By leveraging optimal transport, inference costs are significantly reduced. In the second layer (micro), within regional server clusters, our method assigns specific GPU inference tasks to suitable servers. The quality of server selection not only reduces task completion time but also avoids oscillations and high overhead caused by frequent server state transitions. Through offline training with large-scale samples, the model can quickly solve problems during the online phase.
Specifically, we model the GPU resource scheduling problem as a Markov Decision Process (MDP) to capture the spatial and temporal dynamics of resource allocation. The state space includes current GPU idle rates, inference request queue lengths, network latency matrices, temporal cycle features, and GPU inference task attributes. The action space involves selecting the region or node to handle each inference request. The reward function balances multiple objectives, including average inference latency, load variance, and request drop rate. We employ the Proximal Policy Optimization (PPO) algorithm [36], incorporating decisions from the optimal transport scheme as an additional supervised reward signal to enhance training efficiency. Through interactions with a simulated environment, the system learns optimal multi-time-slot resource allocation strategies. This approach accounts for the temporal dependency of decisions, achieving global optimization across the entire time window rather than myopic optimization within a single time slot.
@@
  • Analysis. We identify the limitations of existing "Reactive" GPU scheduling approaches. These include a lack of foresight and disregard for temporal dependencies, which can significantly increase the completion time of GPU inference tasks and the switching costs of GPU servers. Additionally, compared to reactive systems, our method performs better in scenarios involving cluster failure recovery.
@@
  • Evaluation. We demonstrate through extensive experiments across various scenarios. They include various network topology sizes and multiple metrics. We find that our approach achieves superior response times and completion rates while maintaining balanced loads and lower inference costs compared to existing methods.
We organize the remainder of this paper as follows. In Section §2 and §3, we present our motivation, emphasizing the importance of incorporating the temporal dimension in the GPU scheduling problem. Section §5 provides a detailed description of our methodology, illustrating how reinforcement learning combined with optimal transport is employed to address GPU resource allocation. In Section §6, we conduct extensive experiments, including both simulations and real-world deployments, to demonstrate the effectiveness of our approach. Sections §7 and §8 cover related work, and conclude the paper with conclusions and future directions, respectively.
2 The Need for Temporal-Aware Allocation
@@
2.1 Reactive Scheduling and Its Limitations
@@
Figure 2 illustrates the performance difference between reactive and predictive scheduling when handling periodic traffic peaks. Reactive approach only respond passively after a traffic surge (as seen in the red portion of Figure 2.a), leading to a significant number of requests waiting in queues. Figure 2.b shows the distribution of GPU inference task queuing times during traffic surges, revealing a clear "bimodal pattern" where queuing times are predominantly either long or short, with long queuing times significantly exceeding short ones. In real-world scenarios, this greatly impacts the user experience of LLM services6. In contrast, an ideal predictive approach (the blue dashed line in Figure 2.a) prepares resources 15 minutes in advance through forecasting, resulting in a smoother overall power increase. As shown by the blue bars in Figure 2.b, an ideal predictive approach can effectively reduce the average queuing time to 6 seconds by responding proactively, representing a 53.6% decrease compared to the reactive approach.
Figure 2.c displays the variation in average queuing time throughout the process, showing a clear "staircase effect". The system only begins scaling up after detecting a load increase, and GPUs require 1-3 minutes to transition from cold start to full readiness. During this period, accumulated requests result in an average queuing time of 15.7 seconds. Queuing time gradually decreases as servers scale up, eventually dropping to within 1 second after the GPU request load stabilizes.
@@
2.2 Costs of Ignoring Decision Continuity
@@
These observations reveal a fundamental limitation, reactive methods face a performance upper bound determined by optimal single-timeslot solutions plus unavoidable switching costs. Our analysis in Appendix A.1 proves this upper bound exists and provides conditions for surpassing it through temporal-aware optimization. And we aim to design smoother state transition strategies that reduce switching overhead while maintaining flexibility.
@@
2.3 Allocation Inefficiency Without Coordination
@@
3 Key Challenges
Although introducing temporal-aware mechanisms appears to address some limitations of reactive methods discussed in § 2, their implementation still faces several challenges. These include multi-objective optimization, the difficulty of adapting existing methods to the temporal dimension, and the challenge of optimizing long-term goals.
3.1 Multi-Objective Optimization Complexity
@@
Load Balancing Max 60% Per region Prevents concentration
@@
3.2 How to Optimize for Long-term Benefits?
@@
4 Overview
We propose a scheme called Temporal Optimal Resource scheduling via Two-layer Architecture (TORTA) to address the limitations of reactive scheduling in §2 and the challenges of multi-objective optimization complexity in §3. This scheme systematically addresses the shortcomings of existing methods in handling temporal dependencies, decision continuity, and long-term optimization through temporal dimension modeling and a two-layer optimization architecture.
We remodel GPU resource allocation as a spatio-temporal coupled optimization problem. This considers both spatial heterogeneity (GPU resource distribution and performance differences across geographical regions) and temporal dynamics (load fluctuations, state transition costs, and long-term dependencies). Through this unified spatiotemporal modeling, the system can fully consider the impact of current allocation decisions on future system states when making decisions, thereby avoiding the "short-sighted" problems caused by traditional reactive methods. The optimization objectives include not only immediate task completion time and computational costs, but also minimizing state transition overhead across time slots.
@@ 4 Overview
$\circ$ Macro Regional Allocation (§5.2): This layer handles strategic resource scheduling tasks across regions, focusing on solving geographical distribution imbalances and large-scale load balancing problems. The layer first uses a demand predictor5.2.2) to analyze historical load patterns and temporal features, predicting request distributions for future time slots in each region to provide decision basis for subsequent optimization. The core RL-OT optimizer combines the long-term planning capability of reinforcement learning with the supply-demand matching advantages of optimal transport theory to generate inter-regional allocation strategies that consider temporal continuity.
$\circ$ Micro Server Selection (§5.3): Under the regional allocation constraints determined by the macro layer, this layer focuses on fine-grained task-server matching problems within clusters. The task-server optimization engine comprehensively evaluates hardware compatibility, real-time load status, and server warm-up conditions to compute optimal server allocation schemes for each task, and continuously optimizes initial allocation results through local search algorithms. The state manager coordinates the timing of server state transitions, intelligently deciding the switching timing of each server between different operating states based on load predictions, significantly reducing the high switching costs analyzed in §3 through proactive state management.
5 Design of TORTA
@@ 5.1 Problem Formulation
We formalize the temporal-aware GPU resource allocation problem as a spatiotemporal optimization problem with multiple time slots. Consider a distributed system consisting of $R$ geographical regions, where each region $r\in\{1,2,\ldots,R\}$ contains $S_{r}$ GPU servers. The optimization is performed over a time horizon of $T$ discrete time slots, where each time slot has a duration of $\Delta t$ (typically 5-15 minutes in practice).
@@
The switching cost $C_{switch}^{t}$ directly measures the allocation differences between consecutive timeslots, capturing the overhead of changing resource allocation patterns over time. This cost is crucial for capturing the temporal dependencies identified in our motivation analysis. We define $C_{switch}^{t}=||X_{t}-X_{t-1}||_{F}^{2}$, where $X_{t}$ and $X_{t-1}$ represent the allocation matrices at timeslots $t$ and $t-1$ respectively, and $||\cdot||_{F}$ denotes the Frobenius norm. This formulation directly penalizes dramatic allocation changes between consecutive timeslots, encouraging temporal smoothness while allowing necessary adaptations to changing demand patterns.
@@
$\min_{\pi}\mathbb{E}\Bigg{[}$ $\sum_{t=1}^{T}\bigg{(}\sum_{i=1}^{N_{t}}T_{completion}(i,r,s,t)+\alpha\cdot C_{switch}^{t}+\beta\cdot C_{power}^{t}\bigg{)}\Bigg{]}$ (1)
@@
5.2 Macro-Level Allocation: OT with RL
The macro-level allocation layer addresses the strategic distribution of GPU inference tasks across geographical regions. This layer operates by first establishing a baseline allocation strategy using optimal transport theory, which then serves as a supervised signal for reinforcement learning optimization. The two-stage scheme enables the system to leverage the theoretical optimality guarantees of optimal transport while adapting to complex temporal dynamics through reinforcement learning.
@@ 5.2 Macro-Level Allocation: OT with RL
5.2.1 Optimal Transport for Regional Load Balancing
@@
The cost matrix $C\in\mathbb{R}^{R\times R}$ captures the transportation cost between regions, where $C_{i,j}$ represents the cost of serving a request from region $i$ using resources in region $j$. This cost incorporates network latency, data transfer overhead, and regional operational cost differences: $C_{i,j}=w_{1}\cdot L_{i,j}+w_{2}\cdot BandwidthCost_{i,j}+w_{3}\cdot(OperationalCost_{j}-OperationalCost_{i})$, where $L_{i,j}$ is the network latency between regions $i$ and $j$, and $w_{1}$, $w_{2}$, $w_{3}$ are weighting parameters.
@@
The solution $P^{*}$ provides a baseline allocation strategy where $P^{*}_{i,j}$ represents the optimal mass flow from region $i$ to region $j$. To convert this into actionable routing probabilities, we normalize each row of the allocation matrix: $Prob_{i\rightarrow j}=\frac{P^{*}_{i,j}}{\sum_{k=1}^{R}P^{*}_{i,k}}$. This normalization yields a row-stochastic matrix where $Prob_{i\rightarrow j}$ represents the probability that a task originating from region $i$ should be routed to region $j$ for processing.
The resulting probability distribution
$Prob_{i}=[Prob_{i\rightarrow 1},Prob_{i\rightarrow 2},\ldots,Prob_{i\rightarrow R}]$
for each source region $i$ serves as a reference allocation strategy that balances supply and demand optimally under the given cost structure. However, this static optimal transport solution does not account for temporal dependencies, future load predictions, or the dynamic nature of system states. To address these limitations, we integrate the optimal transport solution as a supervised signal within our reinforcement learning framework.
5.2.2 Reinforcement Learning with OT Supervision

### v2

@@ Abstract
The rapid growth of large language model (LLM) services imposes increasing demands on distributed GPU inference infrastructure. Most existing scheduling systems follow a reactive paradigm, relying solely on the current system state to make decisions, without considering how task demand and resource availability evolve over time. This lack of temporal awareness in reactive approaches leads to inefficient GPU utilization, high task migration overhead, and poor system responsiveness under dynamic workloads. In this work, we identify the fundamental limitations of these instantaneous-state-only scheduling approaches and propose Temporal Optimal Resource scheduling via Two-layer Architecture (TORTA). TORTA introduces a spatiotemporal scheduling framework that captures both long-term workload patterns and short-term execution constraints. It adopts a two-layer design: a macro-level scheduler leverages reinforcement learning and optimal transport to coordinate inter-region task distribution, while a micro-level allocator refines task-to-server assignments within each region to reduce latency and switching costs. Experimental results across multiple network topologies show that TORTA reduces average inference response time by up to 15%, improves load balance by approximately 4-5%, and cuts total operational cost by 10-20% compared to state-of-the-art baseline methods.
Index Terms:
Data Center Network, Reinforcement Learning, LLM Requests Schedule, Optimal Transport
I Introduction
@@
Accelerating GPU inference can be divided into two categories. The first is local acceleration, which involves optimizing the GPU’s operational mode. This includes techniques such as optimizing checkpoint loading, introducing real-time migration and intelligent scheduling strategies, or employing model parallelism and statistical multiplexing to reduce unnecessary overhead [24, 8, 23, 54]. The second category is distributed resource scheduling, which aims to improve GPU resource utilization and prevent GPU server idleness [49]. Some approaches focus on allocating suitable GPU resources to different tasks to achieve the fastest inference, often with additional optimization objectives such as minimizing communication costs and power consumption [40, 16].
@@
The complexity of this scheduling problem is often very high. Finding an optimal scheduling solution typically requires significant computational time and increases dramatically with network scale expansion, severely impacting task completion time. Existing approaches usually seek approximately optimal solutions within acceptable time constraints, and these methods have achieved good results to a certain extent. For example, MLServerless [8] employs an offline-plus-online strategy, where the offline phase invests substantial effort in learning solution strategies, enabling the system to quickly derive good solutions during the online phase to achieve optimal inference speed. Other approaches [16, 40] model this as a Mixed Integer Linear Programming (MILP) or Integer Quadratic Programming (IQP) problem and accelerate solution speed through solution space pruning. However, most of these approaches are ”Reactive”, making immediate decisions based on currently observed states (such as cluster interference levels, resource utilization rates, and energy prices), essentially finding optimal solutions within individual time slots. This paradigm inherently ignores the temporal perspective and continuity considerations in decision-making.
@@
Given these challenges, our approach addresses GPU resource allocation by integrating reinforcement learning with an optimal transport [41] perspective, particularly incorporating the temporal dimension. Unlike traditional methods, our strategy—termed Temporal Optimal Resource scheduling via Two-layer Architecture (TORTA)—divides the allocation process into two hierarchical layers: macro and micro. In the first layer (macro), TORTA assigns GPU inference tasks to nodes in each region, primarily to mitigate the imbalance between GPU resources and task distribution. By leveraging optimal transport, inference costs are significantly reduced. In the second layer (micro), within regional server clusters, TORTA assigns specific GPU inference tasks to suitable servers. The quality of server selection not only reduces task completion time but also avoids oscillations and high overhead caused by frequent server state transitions. Through offline training with large-scale samples, the model can quickly solve problems during the online phase.
Specifically, we model the GPU resource scheduling problem as a Markov Decision Process (MDP) with states encompassing GPU utilization, queue lengths, network latency, and temporal features. The action space involves regional allocation decisions, while the reward function balances response time, load variance, and operational cost. We employ Proximal Policy Optimization (PPO) [35] with optimal transport decisions as supervised signals, enabling the system to learn multi-time-slot allocation strategies that achieve global optimization rather than myopic single-slot decisions.
@@
  • Analysis. We identify the limitations of existing ”Reactive” GPU scheduling approaches. These include a lack of foresight and disregard for temporal dependencies, which can significantly increase the completion time of GPU inference tasks and the switching costs of GPU servers. Additionally, we establish performance upper bound for these methods and derive explicit conditions under which temporal-aware optimization can provably outperform this bound.
@@
  • Evaluation. We demonstrate through extensive experiments across various scenarios. They include four different network topologies and multiple performance metrics. We find that our approach achieves superior response times (16.39s-19.31s vs. 18.72s-24.39s for baselines) and load balance coefficients (0.74-0.76 vs. 0.68-0.73 for baselines) while maintaining lower operational costs with power cost reductions of 10.7K-14.1K compared to existing methods.
We organize the remainder of this paper as follows. In Section II and III, we present our motivation, emphasizing the importance of incorporating the temporal dimension in the GPU scheduling problem. Section V provides a detailed description of our methodology, illustrating how reinforcement learning combined with optimal transport is employed to address GPU resource allocation. In Section VI, we conduct extensive experiments to demonstrate the effectiveness of our approach. Sections VII and VIII cover related work, and conclude the paper with conclusions and future directions, respectively. Detailed theoretical proofs for performance guarantees, comprehensive network training procedures, and additional experimental results are provided in the Appendix as supplemental material.
II Motivation:
The Need for Temporal-Aware Allocation
@@ II Motivation:
II-A Reactive Scheduling and Its Limitations
@@
Figure 2 illustrates the performance difference between reactive and predictive scheduling when handling periodic traffic peaks. Reactive approach only respond passively after a traffic surge (as seen in the red portion of Figure 2.a), leading to a significant number of requests waiting in queues. Figure 2.b shows the distribution of GPU inference task queuing times during traffic surges, revealing a clear ”bimodal pattern” where queuing times are predominantly either long or short, with long queuing times significantly exceeding short ones. In real-world scenarios, this greatly impacts the user experience of LLM services. In contrast, an ideal predictive approach (the blue dashed line in Figure 2.a) prepares resources 15 minutes in advance through forecasting, resulting in a smoother overall power increase.
Figure 2.c displays the variation in average queuing time throughout the process, showing a clear ”staircase effect”. The system only begins scaling up after detecting a load increase, and GPUs require 1-3 minutes to transition from cold start to full readiness. During this period, accumulated requests result in an average queuing time of 15.7 seconds. Queuing time gradually decreases as servers scale up, eventually dropping to within 1 second after the GPU request load stabilizes.
@@
II-B Costs of Ignoring Decision Continuity
@@
These observations reveal a fundamental limitation, reactive methods face a performance upper bound determined by optimal single-timeslot solutions plus unavoidable switching costs. Our analysis in Appendix proves this upper bound exists and provides conditions for surpassing it through temporal-aware optimization. And we aim to design smoother state transition strategies that reduce switching overhead while maintaining flexibility.
@@
II-C Allocation Inefficiency Without Coordination
@@
III Key Challenges
Although introducing temporal-aware mechanisms appears to address some limitations of reactive methods discussed in § II, their implementation still faces several challenges.
III-A Multi-Objective Optimization Complexity
@@
Load Balancing Max 80% Per region Prevents concentration
@@
III-B How to Optimize for Long-term Benefits?
@@
IV Torta Overview
We propose a scheme called Temporal Optimal Resource scheduling via Two-layer Architecture (TORTA) to address the limitations of reactive scheduling in §II and the challenges of multi-objective optimization complexity in §III. This scheme systematically addresses the shortcomings of existing methods in handling temporal dependencies, decision continuity, and long-term optimization through temporal dimension modeling and a two-layer optimization architecture.
We remodel GPU resource allocation as a spatio-temporal coupled optimization problem. This considers both spatial heterogeneity (GPU resource distribution and performance differences across geographical regions) and temporal dynamics (load fluctuations, state transition costs, and long-term dependencies). Through this unified spatiotemporal modeling, the system can fully consider the impact of current allocation decisions on future system states when making decisions, thereby avoiding the ”short-sighted” problems caused by traditional reactive methods. The optimization objectives include not only immediate task completion time and computational costs, but also minimizing state transition overhead across time slots.
@@ IV Torta Overview
$\circ$ Macro Regional Allocation (§V-B): This layer handles strategic resource scheduling tasks across regions, focusing on solving geographical distribution imbalances and large-scale load balancing problems. The layer first uses a demand predictorV-B2) to analyze historical load patterns and temporal features, predicting request distributions for future time slots in each region to provide decision basis for subsequent optimization. The core RL-OT optimizer combines the long-term planning capability of reinforcement learning with the supply-demand matching advantages of optimal transport theory to generate inter-regional allocation strategies that consider temporal continuity.
$\circ$ Micro Server Selection (§V-C): Under the regional allocation constraints determined by the macro layer, this layer focuses on fine-grained task-server matching problems within clusters. The task-server optimization engine comprehensively evaluates hardware compatibility, real-time load status, and server warm-up conditions to compute optimal server allocation schemes for each task, and continuously optimizes initial allocation results through local search algorithms. The state manager coordinates the timing of server state transitions, intelligently deciding the switching timing of each server between different operating states based on load predictions, significantly reducing the high switching costs analyzed in §3 through proactive state management.
V Design of TORTA
@@ V-A Problem Formulation
We formalize the temporal-aware GPU resource allocation problem as a spatiotemporal optimization problem with multiple time slots. Consider a distributed system consisting of $R$ geographical regions, where each region $r\in\{1,2,\ldots,R\}$ contains $S_{r}$ GPU servers. The optimization is performed over a time horizon of $T$ discrete time slots, where each time slot has a duration of $\Delta t$.
@@
[…]
@@
$\min_{\pi}\mathbb{E}\Bigg{[}$ $\sum_{t=1}^{T}\bigg{(}\sum_{i=1}^{N_{t}}T_{completion}(i,r,s,t)$ (1)
$\quad+\alpha\cdot C_{switch}^{t}+\beta\cdot C_{power}^{t}\bigg{)}\Bigg{]}$
@@
V-B Macro-Level Allocation: OT with RL
The macro-level allocation layer addresses the strategic distribution of GPU inference tasks across geographical regions. This layer operates by first establishing an upper-bound allocation strategy using optimal transport theory (as proved in Appendix), which then serves as a supervised signal for reinforcement learning optimization. The two-stage scheme enables the system to leverage the theoretical optimality guarantees of optimal transport while adapting to complex temporal dynamics through reinforcement learning.
@@ V-B Macro-Level Allocation: OT with RL
V-B1 Optimal Transport for Regional Load Balancing
@@
The cost matrix $C\in\mathbb{R}^{R\times R}$ captures the cost between regions, where $C_{i,j}$ represents the cost of serving a request from region $i$ using resources in region $j$. This cost primarily incorporates regional power cost differences and minimal network overhead: $C_{i,j}=w_{1}\cdot(PowerCost_{j})+w_{2}\cdot(L_{i,j}+BandwidthCost_{i,j})$, where $PowerCost_{j}$ represents the electricity cost in region $j$, $L_{i,j}$ is the network latency between regions $i$ and $j$, $BandwidthCost_{i,j}$ accounts for data transfer overhead, and $w_{1}\gg w_{2}$ reflects the dominance of power costs over network costs in the optimization objective.
@@
The solution $P^{*}$ provides an upper-bound allocation strategy where $P^{*}_{i,j}$ represents the optimal mass flow from region $i$ to region $j$. To convert this into actionable routing probabilities, we normalize each row of the allocation matrix: $Prob_{i\rightarrow j}=\frac{P^{*}_{i,j}}{\sum_{k=1}^{R}P^{*}_{i,k}}$. This normalization yields a row-stochastic matrix where $Prob_{i\rightarrow j}$ represents the probability that a task originating from region $i$ should be routed to region $j$ for processing.
The resulting probability distribution $Prob_{i}=[Prob_{i\rightarrow 1},Prob_{i\rightarrow 2},\ldots,Prob_{i\rightarrow R}]$ for each source region $i$ serves as a reference allocation strategy that balances supply and demand optimally under the given cost structure. However, this static optimal transport solution does not account for temporal dependencies, future load predictions, or the dynamic nature of system states. To address these limitations, we integrate the optimal transport solution as a supervised signal within our reinforcement learning framework.
V-B2 Reinforcement Learning with OT Supervision